can you help me with this last problem; solve using the 5 step problem process


a football is about 50 yards wide by 120 yards long. the border of uniform width around the entire field has an area of about 1800 square yards. how wide is the border around the field?

thank you:)

Found 2 solutions by richwmiller, Nate:
Answer by richwmiller(17219)

sorry but there is no football that big.

Answer by Nate(3500)

At first glance, this can seem quite a complex geometrical problem. I always find that drawing a picture very much simplifies the mathematical reasoning. A rectangle has length 120 and width 50. A second rectangle surrounds the first with uniform width. Therefore, the length of the larger rectangle is 120 + 2x and the width is 50 + 2x
The area of the large rectangle = length*width = (120 + 2x)(50 + 2x)
The area of the smaller rectangle = 120*50 = 6000
The difference of these areas will result in the area of the border which runs along the smaller rectangle, or football field.
1800 = (120 + 2x)(50 + 2x) - 6000
You can solve the rest.